OpenStudy (anonymous):

tim has $20.00 to spend on 6 pens. After buying them, he had $8.00. How much did each pen cost?

OpenStudy (anonymous):

subtract 8 from 20 then divide by 6

OpenStudy (anonymous):

Initial amount of money he had = $20 Balance after buying the pens = $8 So amount of money spent on the pens = 20 - 8 = $12 No. of pens Tim bought = 6 pens So amount spent on each pen = 12/6 = $2

OpenStudy (anonymous):

thank you everyone

OpenStudy (anonymous):

Those pens are expensive